# GazeNoter: Co-Piloted AR Note-Taking via Gaze Selection of LLM Suggestions to Match Users’ Intentions

### v1

@@ 3.2.1. Context Keyword Selection
The system captures a voice segment, subsequently displaying context keywords via the LLM keyword extraction in real-time from the latest sentence. As shown in Figure 2 (a), these context keywords appear on the top section of the AR display in a 1 $\times$ 4 layout, at most 4 keywords from the most recent sentence added in the last of the queue clearly underlined. Users can also view previous context keywords using the previous/next arrow buttons. Users select context keywords relevant to the content they want to record as a note (e.g., “city” in Figure 2 (b)), and the selected keywords are showcased on the right-hand side of the AR display. The selected keywords serve as quick notes, allowing users to express their thoughts and serve as reminders in a short period without recording complete sentences (Chen et al., 2021). Notably, only context keywords are displayed on the top section of the AR headset when users touch the button of the ring if no keyword is selected to prevent the AR contents from interfering with real-world activities.
@@ 3.2.3. Derivative Keyword Selection
Customized keyword selection involving users’ habits and intentions might generate notes slightly beyond the context. However, to indeed accomplish beyond-context notes, keywords other than context keywords are required. In our brains, we usually think of something or some ideas when we are inspired by the things we see or hear, and we combine those things with our background knowledge, experiences and even personality to generate the thoughts. Therefore, when taking a beyond-context note and cannot find proper context keywords, users can opt to double-click on the most relevant context keyword (e.g., “rallies” in Figure 2 (d)) to generate derivative keywords via the LLM keyword derivation. The original keyword and 4 derivative keywords, a total of 5 items, are presented in a circular layout at the center of the AR display. Users can select derivative keywords that best match their intentions (e.g., “sign” in Figure 2 (e)), view more derivative keywords via the previous/next arrow buttons, or double-click on a derivative keyword close to their intentions to further obtain more relevant derivative keywords generated by the LLM. Derivative keyword selection could take more time but obtain more precise notes matching users’ intentions for beyond-context notes. Therefore, it is a trade-off for users.
@@ 3.2.5. Note Refinement
In general, users primarily select context keywords and may also select a customized keyword(s), and then select a candidate sentence as a note or directly record selected keywords as a note. Furthermore, they can select only context (or also customized) keywords to record these as a quick note in just two or three selection steps, which means that a quick note is defined as one recorded only from (a) (b) or (a) (b) (c) steps in Figure 2 in GazeNoter. Performing a few keyword and sentence selection steps to take notes accomplishes the real-time system (DC1) and distraction and cognitive load (DC3) design considerations. Using context keyword selection, customized keyword selection, and candidate sentence selection, the system matches users’ intentions for within-context notes. Advanced derivative selection is used only for beyond-context notes, and by incorporating this feature, the system fulfills the matching users’ intentions design consideration (DC2).
@@ 3.3. AR Gaze Selection
For input, compared with gesture, controller and voice input, gaze selection is a more swift and subtle input approach, and the eye-tracking system is commonly built in off-the-shelf headsets, e.g., Microsoft Hololens 2, Meta Quest Pro, Apple Vision Pro and HTC Vive Pro Eye. However, although several gaze confirmation methods are proposed, e.g., dwelling and gaze gestures, pressing a button or tapping or clicking gestures are still more practical and commonly-used gaze confirmation approaches currently. Therefore, we designed a ring, consisting of a button, an infrared (IR) sensor and spiral tracks, worn on the index finger for confirmation, as shown in Figure 4. Users press the button with their thumb to confirm the gaze-selected item. Using a ring instead of a controller, users can even perform the confirmation using the ring when putting their hand in a pocket during discussions or conversations, accomplishing better subtlety and mobility. This cannot be achieved by input and confirmation methods using gestures since current gesture tracking on AR headsets is mainly camera-based, suffering from the occlusion problem. Certainly, the ring can be replaced by any advanced confirmation approaches in the future. The combination of AR gaze selection and ring confirmation fulfills not only swift (DC1) and subtle (DC4) but also low-distraction input (DC3) in the design considerations.
@@
Besides gaze selection, rapidly turning the display on and off to achieve rapid switching between virtual and real worlds is another merit of AR headsets. Therefore, the IR sensor (QTR-1RC Reflectance Sensor) is equipped on the button of the ring, and a small case with a small hole is above the IR sensor. When the thumb touches the button and covers the small hole, the content is displayed on the AR headset, and the thumb can further press the button with the mechanical “click” feedback for confirmation of gaze selection. Double-click is also used for the alternative confirmation, which means two consecutive clicks within 500ms. Once the thumb leaves the button, the display is turned off, and users can clearly see the real world. The touch gesture via the IR sensor enables users to rapidly switch between AR content and real-world activities. Furthermore, whenever the thumb touches the button, the AR content in the designed layout would be realigned to the current headset direction. Therefore, users can dynamically adjust the AR content position by re-touching the button when the real-world target is in the blank center of the headset in our AR layout. In addition, as long as the thumb stays on the button, the AR layout remains fixed in the physical space. Thus, users can select items close to the edges of the field of view (FoV) of the AR display easily by slightly moving the head. Notably, all AR items in the layout are within the AR display, so users can only move their eyes to achieve AR gaze selection, and the head movement is only auxiliary.
@@
The proposed AR gaze selection with the ring achieves swift and subtle gaze selection, rapid switching across reality, AR content position adjustment, and conventional ring form factor. The button and IR sensor are controlled by a NodeMCU-32S microcontroller. The weight of the 3D-printed ring and the sensors is 4.9g. The lightweight ring and the AR headset, which could be envisioned to be as small as normal glasses in the future, in GazeNoter accomplish the mobility design consideration (DC5).
@@ 3.4.1. Context Keyword Extraction
In the beginning, the audio speech is received by the headset and is then converted into text using the pre-trained model, Whisper, provided by OpenAI (size of the model: base, English-only model). Every 4 seconds we record a voice segment as the input to the Whisper model. If there is a pause of over 1 second in records, it is treated as the start of a new sentence in the transcription. The output transcript is used as input for the LLM to generate context keywords. We observed that a sentence or part of a long sentence with a comma is usually said in 4 seconds, so this can generate context keywords of the sentence in real time. The average delay for context keyword generation is about 4.29 seconds after users hear the sentence. In addition to the current sentence, previous sentences in the transcripts are also used as input for the LLM. These provide context for the LLM to generate more precise context keywords of the sentence. The LLM was prompted to extract at most 4 keywords from each sentence. This number range was chosen based on a pilot study to avoid redundancy and ensure that essential keywords are included.
@@ 3.4.2. Derivative Keyword Derivation
For beyond-context notes, users select the context keyword most relevant to their intentions and use it as input for the LLM to generate 4 derivative keywords, which are relevant to and derived from the context keyword. This is like using the LLM as a brain to derive or associate relevant keywords from/with the context keyword for users. 4 derivative keywords consist of 2 types, including 2 derivative keywords contextually associated with the original keyword, which are relevant to both the context and the original context keyword, and 2 derivative keywords exclusively related to the original keyword, which are not relevant to the context but only relevant to the original context keyword. The prompts for both types instruct the LLM to generate related words while explicitly prohibiting the generation of words that overlap with the context keywords currently displayed to the users or have the same lemma of a word. This maximizes the number of new keywords present to the users. The prompt for derivative keywords contextually associated with the original keyword includes the preceding 15 sentences as the context for the LLM. We observed that 15 sentences can provide adequate context without including older topics unrelated to the current context based on a pilot study. When users intend to further obtain more proper derivative keywords by double-clicking on a derivative keyword close to their intentions, the derivative keyword is used as input for LLM to generate more derivative keywords. That original derivative keyword and the other 4 derivative keywords exclusively related to that original derivative keyword are shown in the circular layout. The average delay of derivative keyword derivation is about 1.41 seconds after double-clicking a keyword.
@@ 3.4.3. Candidate Sentence Organization
Once a keyword is selected, the system automatically uses all selected keywords as input of the LLM to generate 3 candidate sentences. We prompt the LLM to ensure that each sentence either contains all selected keywords or shares the same lemma of words with the selected keywords. We also limit each candidate sentence to 10 words for ease of reading. Furthermore, if any WH words or a question mark from customized keywords is selected, the candidate sentences must be in the form of questions. The average delay of candidate sentence generation is about 2.89 seconds.
@@ 4. User Study 1: Formal Speech
To understand how users leverage the GazeNoter system, verify its usability, and compare it with other note-taking methods in a static speech scenario, we conducted this user study.
@@ 4.1. Participants and Apparatus
An Oculus Quest Pro headset with a built-in eye-tracking system was used. 5 ring sizes, ranging from 18mm to 22mm in diameter, were prepared. The ring for gaze confirmation was worn on the index finger. A smartphone, iPhone 8 plus (5.5-inch display), was used for comparison. We recruited 12 participants (5 female) aged 19-27 (mean: 24.16) were recruited.
@@ 4.2. Task and Procedure
To simulate attending speeches, 6 TED Talks videos with similar difficulty and length, about 10 minutes, were used. Three note-taking methods were compared, including smartphone text entry (T), smartphone version GazeNoter (S), and the proposed AR GazeNoter (G). An iPhone built-in note-taking app, Notes, was used in the smartphone text entry (T), which was used as a baseline. For the smartphone version GazeNoter (S), we implemented the proposed system on the smartphone. All contents were shown on the smartphone screen, and the layout was similar to that on the AR display but was slightly adjusted to fit in the commonly-used portrait mode of the smartphone (Figure 5). The font size was the same as in the built-in note-taking app in (T). Since the blank center was not needed on the smartphone, the circular layout for derivative keywords was not used. Touch input was used to select items on the smartphone via tap and double-tap gestures. By comparing (T) and (S) on the smartphone, we could understand whether our users-in-the-loop LLM system outperformed the current manual text entry note-taking methods. By comparing (S) and (G), we could observe the distraction, cognitive load subtlety, and social acceptance of GazeNoter on the smartphone and AR headset. (S) and (G) are both methods proposed in this paper but on different media. Notably, we also intended to observe whether the notes from the proposed user-in-the-loop LLM system outperformed automatically generated notes by the LLM. Therefore, notes were automatically generated by using the whole speech transcript for the LLM after the speech in (G). This was an additional automatic note-taking method (A) as another baseline although no user input was in this method. The generated note number in (A) was the same as in (G) by each participant. Thus, this additional comparison was only between (A) and (G).
@@
Participants sat on a chair and listened to speeches. After the experiment and GazeNoter were introduced, an eye-tracking calibration on the AR headset was performed for each participant, and a proper ring was worn. The speech videos were played on the AR display in a fixed space in the real world in (G) to simulate real speeches and on a monitor (32-inch display) on the desk in front of them in (T) and (S). To prevent the noise in the environment from interfering with the performance, our system directly received the audio source from the computer for speech-to-text conversion when playing the videos. A training session was performed after the introduction. One video was used in the training for all participants. We encouraged them to explore every feature of GazeNoter, and asked them to think about their quick note-taking strategies for GazeNoter based on their experiences in the training session. During the experiment, they watched a TED Talks video in each method and performed note-taking for questions in the Q&A sessions or important points and ideas they wanted to record. After the speeches, they could review the notes, see the transcripts of the notes, or even perform refinement in a 30-second time window, as a small time window between a speech and the Q&A session. Certainly, these could be performed during the speeches.

### v2

@@ 3.2.1. Context Keyword Selection
The system captures a voice segment, subsequently displaying context keywords via the LLM keyword extraction in real-time from the latest sentence. As shown in Figure 3 (a), these context keywords appear on the top section of the AR display in a 1 $\times$ 4 layout, at most 4 keywords from the most recent sentence added in the last of the queue clearly underlined. Users can also view previous context keywords using the previous/next arrow buttons. Users select one or multiple context keyword(s) relevant to the content they want to record as a note (e.g., “city” in Figure 3 (b)), and the selected keywords are showcased on the right-hand side of the AR display. The selected keywords serve as quick notes, allowing users to express their thoughts and serve as reminders in a short period without recording complete sentences (Chen et al., 2021). Notably, only context keywords are displayed on the top section of the AR headset when users touch the button of the ring if no keyword is selected to prevent the AR contents from interfering with real-world activities.
@@ 3.2.3. Derivative Keyword Selection
Customized keyword selection involving users’ habits and intentions might generate notes slightly beyond the context. However, to indeed accomplish beyond-context notes, keywords other than context keywords are required. In our brains, we usually think of something or some ideas when we are inspired by the things we see or hear, and we combine those things with our background knowledge, experiences and even personality to generate the thoughts. Therefore, when taking a beyond-context note and cannot find proper context keywords, users can opt to double-click on the most relevant context keyword (e.g., “rallies” in Figure 3 (d)) to generate derivative keywords via the LLM keyword derivation. The original keyword and 4 derivative keywords, a total of 5 items, are presented in a circular layout at the center of the AR display. Users can select derivative keywords that best match their intentions (e.g., “sign” in Figure 3 (e)), view more derivative keywords via the previous/next arrow buttons, or double-click on a derivative keyword close to their intentions to further obtain more relevant derivative keywords generated by the LLM. Although the concepts of personal experiences and habits are used in customized keywords, these were not incorporated into derivative keywords. However, since users can further generate more relevant derivative keywords based on selected derivative keywords, even across multiple iterations if desired, this iterative process allows the newly generated derivative keywords to progressively narrow down and align more closely with users’ intentions. Derivative keyword selection could take more time but obtain more precise notes matching users’ intentions for beyond-context notes. Therefore, it is a trade-off between time and precision for users.
@@ 3.2.5. Note Refinement
(Left) Users could review the notes by pressing the red ”Notes” button, always on the display. (Right) When further selecting a note, they can see the transcripts of the selected keywords of the note. The note and transcripts as well as the keywords and candidate sentences are shown to remind users of the context at the time.
In general, users primarily select context keywords and may also select a customized keyword(s), and then select a candidate sentence as a note or directly record selected keywords as a note. Furthermore, they can select only context (or also customized) keywords to record these as a quick note in just two or three selection steps, which means that a quick note is defined as one recorded only from (a) (b) or (a) (b) (c) steps in Figure 3 in GazeNoter. Performing a few keyword and sentence selection steps to take notes accomplishes the real-time system (DC1) and distraction and cognitive load (DC3) design considerations. Using context keyword selection, customized keyword selection, and candidate sentence selection, the system matches users’ intentions for within-context notes. Advanced derivative selection is used only for beyond-context notes, and by incorporating this feature, the system fulfills the matching users’ intentions design consideration (DC2). Although taking notes while staying engaged in the ongoing activity could more or less cause distraction, GazeNoter allows users to freely employ their strategies to take notes based on the density of information in the speech-based activity. Instead of complete sentences, quick notes with only context (or also customized) keywords could be recorded in high-density information parts of the speech-based activity to avoid missing content. Furthermore, two types of historical information are preserved, allowing users to review them in less stressed parts. First, all previous context keywords can be viewed using the previous/next arrow buttons. Second, in note refinement, all selected keywords and their transcripts are also recorded. Therefore, users can add notes even if they have missed some content. Notably, since notes automatically generated by current LLMs could be comprehensive without any miss and could be integrated in our system in the future, our goal is to precisely generate notes match users’ intentions for both within-context and beyond-context notes.
@@ 3.3. AR Gaze Selection
(a) The hardware structure of the ring, consisting of a button, an infrared (IR) sensor on top of the button and spiral tracks, worn on the index finger for confirmation. (b) The button can be withdrawn on the back of the finger, preventing interference with users. (c) The button is extended through the spiral track for input. (d) The ring can be used subtly, such as in a pocket.
@@
For input, compared with gesture, controller and voice input, gaze selection is a more swift and subtle input approach, and the eye-tracking system is commonly built in off-the-shelf headsets. Although several gaze confirmation methods are proposed, e.g., dwelling and gaze gestures, the combination of gaze selection and gesture or ring confirmation has become popular and practical on AR headsets, such as Microsoft Hololens 2, Meta Quest Pro, Meta Orion and Apple Vision Pro. Therefore, we designed a ring, consisting of a button, an infrared (IR) sensor and spiral tracks, worn on the index finger for confirmation, as shown in Figure 5. Users press the button with their thumb to confirm the gaze-selected item. Users can even perform the confirmation using the ring when putting their hand in a pocket during discussions or conversations, accomplishing better subtlety and mobility. This cannot be achieved by gesture input and confirmation due to the occlusion problem in gesture tracking. The ring can be substituted by any advanced confirmation approaches in the future, such as picoRing (Takahashi et al., 2024). Besides, since 2D input on a small ring could suffer from a large control-display (CD) ratio issue, we chose the common combination of gaze selection and ring confirmation. This combination fulfills not only swift (DC1) and subtle (DC4) but also low-distraction input (DC3) in the design considerations.
@@
Besides gaze selection, rapidly turning the display on and off to achieve rapid switching between virtual and real worlds is another merit of AR headsets. An IR sensor (QTR-1RC Reflectance Sensor) is equipped on the button of the ring, and a small case with a small hole is above the IR sensor. When the thumb touches the button and covers the hole, the AR content is displayed, and the thumb can further press the button with the mechanical “click” feedback for confirmation. Double-click is also used for the alternative confirmation, which means two consecutive clicks within 500ms. Once the thumb leaves the button, the display is turned off. Therefore, The touch gesture via the IR sensor enables users to rapidly switch between AR content and real-world activities. Notably, a solid background for the text of the AR content is used to prevent the font color from blending with the dynamically changing real-world colors. Although a transparent background could minimize distraction, a solid background with the rapid switching design via the touch gesture reduces distraction and achieves a trade-off between distraction and the clarity and readability of the AR content. Despite using a solid background, the AR layout leaves the center blank, as in  (Cai et al., 2023; Janaka et al., 2022). Therefore, users can still see through the display. With the rapid switching design on the ring, users can spend most of their time without AR content. In addition, whenever the thumb touches the button, the AR content in the designed layout is anchored in the physical space relative to the headset direction at the initial moment of touching the ring. Users can dynamically adjust the AR content position by re-touching the button when the real-world target is in the blank center of the AR layout. As long as the thumb stays on the button, the AR layout remains fixed in the physical space. Thus, users can select items close to the edges of the field of view (FoV) of the AR display easily by slightly moving the head. Notably, all AR items in the layout are within the AR display, so users can only move their eyes to achieve AR gaze selection, and the head movement is only auxiliary.
@@
Although gaze selection with ring confirmation has been explored, the two innovative features of the ring well-suited to our system include: (1) rapid switching between AR content and real-world activities via an IR sensor in a single touch-and-press step to reduce distraction and mitigate the blocking issue, and (2) AR content position adjustment achieved by touching the ring for mobile, real-time conditions. The proposed AR gaze selection with the ring achieves swift and subtle gaze selection, rapid switching across reality, AR content position adjustment, and conventional ring form factor. While the performance of the input method was not evaluated since it is not the primary contribution of this paper, we still verified and discussed the merits and usability of the designs in both static and mobile note-taking scenarios in the following studies. The button and IR sensor are controlled by a NodeMCU-32S microcontroller. The weight of the 3D-printed ring and the sensors is 4.9g. The lightweight ring and the AR headset, which could be envisioned to be as small as normal glasses in the future, in GazeNoter accomplish the mobility design consideration (DC5).
@@ 3.4.1. Context Keyword Extraction
In the beginning, the audio speech is received by the headset and is then converted into text using the pre-trained model, Whisper, provided by OpenAI (size of the model: base, English-only model). Every 4 seconds we record a voice segment as the input to the Whisper model. If there is a pause of over 1 second in records, it is treated as the start of a new sentence in the transcription. The output transcript is used as input for the LLM to generate context keywords. We observed that a sentence or part of a long sentence with a comma is usually said in 4 seconds, so this can generate context keywords of the sentence in real time. The average delay for context keyword generation is about 4.29 seconds after users hear the sentence. In addition to the current sentence, previous sentences in the transcripts are also used as input for the LLM. These provide context for the LLM to generate more precise context keywords of the sentence. The LLM was prompted to extract at most 4 keywords from each sentence. This number range was chosen based on a pilot study to avoid redundancy and ensure that essential keywords are included. We originally designed a 2 $\times$ 5 layout to display 10 context keywords, enabling users to view context keywords from the two most recent simultaneously. However, considering the guideline that the item size for gaze input should exceed $2^{\circ}$ in (Kim et al., 2022) and the precision of the eye-tracking system in off-the-shelves AR headsets, a 1 $\times$ 4 layout was finally adopted. This design ensures that the item size is much larger than $2^{\circ}$ and the context keywords are placed sparser.
@@ 3.4.2. Derivative Keyword Derivation
For beyond-context notes, users select the context keyword most relevant to their intentions and use it as input for the LLM to generate 4 derivative keywords, which are relevant to and derived from the context keyword. This is like using the LLM as a brain to derive or associate relevant keywords from/with the context keyword for users. 4 derivative keywords consist of 2 types, including 2 derivative keywords contextually associated with the original keyword, which are relevant to both the context and the original context keyword, and 2 derivative keywords exclusively related to the original keyword, which are not relevant to the context but only relevant to the original context keyword. The prompts for both types instruct the LLM to generate related words while prohibiting those overlapping with the displayed context keywords or having the same lemma. This maximizes the number of new keywords present to the users. The prompt for derivative keywords contextually associated with the original keyword includes the preceding 15 sentences as the context for the LLM. We observed that 15 sentences can provide adequate context without including older topics unrelated to the current context based on a pilot study. The prompt rule ensures that derivative keywords are related to the original keyword but remain distinct by avoiding overlap with current and past keywords and using different lemmas. When users intend to further obtain more proper derivative keywords by double-clicking on a derivative keyword close to their intentions, the derivative keyword is used as input for LLM to generate more derivative keywords. That original derivative keyword and the other 4 derivative keywords exclusively related to that original derivative keyword are shown in the circular layout. We initially aimed to allocate 6 or even 9 derivative keywords, including the original keyword, in a circular layout, but 5 items in a circular layout enable easier and more accurate derivative keyword selection. The average delay of derivative keyword derivation is about 1.41 seconds after double-clicking a keyword.
@@ 3.4.3. Candidate Sentence Organization
Once a keyword is selected, the system automatically uses all selected keywords as input of the LLM to generate 3 candidate sentences. We prompt the LLM to ensure that each sentence either contains all selected keywords or shares the same lemma of words with the selected keywords. Furthermore, the context of the previous transcripts is included. We also limit each candidate sentence to 10 words for ease of reading. Furthermore, if any WH words or a question mark from customized keywords is selected, the candidate sentences must be in the form of questions. The average delay of candidate sentence generation is about 2.89 seconds. The layout of 3 candidate sentences on the bottom section of AR display is based on the design in (Choi et al., 2022), aiming to make it more comfortable for users to read the sentences, which require longer reading time.
@@ 4. User Study 1: Formal Speech
To understand how users leverage the GazeNoter system, verify its usability, and compare it with other note-taking methods in a static speech scenario, we conducted this user study. Notably, instead of assessing user comprehension or memory retention in note-taking, our primary goal was to evaluate the practical utility and gather subjective feedback on the note-taking experience with GazeNoter. Therefore, we followed the established methodologies for evaluating novel interaction techniques and systems, especially for LLM interactions  (Jiang et al., 2023; Suh et al., 2023) and note-taking  (Chandrasegaran et al., 2019; Fang et al., 2021; Haliburton et al., 2023), in this user study.
@@ 4.1. Participants and Apparatus
An Oculus Quest Pro headset with a built-in eye-tracking system was used. 5 ring sizes, ranging from 18mm to 22mm in diameter, were prepared. The ring for gaze confirmation was worn on the index finger. A smartphone, iPhone 8 plus (5.5-inch display), was used for comparison. 12 participants (5 female) aged 19-27 (mean: 24.16) were recruited. They were compensated with 10 USD for their time.
@@
Study 1 setup of GazeNoter on the AR headset (G) (left) and on the smartphone (S) (middle) with its layout on the smartphone (right).
@@ 4.2.1. Task
To simulate attending speeches, 6 TED Talks videos with similar difficulty and length, about 10 minutes, were used. The TED Talks are similar to lecture scenarios that were extensively studied in prior research (Mosleh et al., 2016; Piolat et al., 2005) but under more time constraints for note-taking. Three note-taking methods were compared, including smartphone text entry (T), smartphone version GazeNoter (S), and the proposed AR GazeNoter (G). An iPhone built-in note-taking app, Notes, was used in the smartphone text entry (T), which was used as a baseline. For the smartphone version GazeNoter (S), we implemented the proposed system on the smartphone. All contents were shown on the smartphone screen, and the layout was similar to that on the AR display but was slightly adjusted to fit in the commonly-used portrait mode of the smartphone (Figure 6). The font size was the same as in the built-in note-taking app in (T). Since the blank center was not needed on the smartphone, the circular layout for derivative keywords was not used. Touch input was used to select items on the smartphone via tap and double-tap gestures. By comparing (T) and (S) on the smartphone, we could understand whether our users-in-the-loop LLM system outperformed the current manual text entry note-taking methods. By comparing (S) and (G), we could observe the distraction, cognitive load subtlety, and social acceptance of GazeNoter on the smartphone and AR headset. (S) and (G) are both methods proposed in this paper but on different media. Notably, we also intended to observe whether the notes from the proposed user-in-the-loop LLM system outperformed automatically generated notes by the LLM. Therefore, notes were automatically generated by using the whole speech transcript for the LLM after the speech in (G). This was an additional automatic note-taking method (A) as another baseline although no user input was in this method. The generated note number in (A) was the same as in (G) by each participant. Thus, this additional comparison was only between (A) and (G).
4.2.2. Procedure
@@
Participants sat on a chair and listened to speeches. After the experiment and GazeNoter were introduced, an eye-tracking calibration on the AR headset was performed, and a proper ring was worn. The speech videos were played on the AR display in a fixed space in the real world in (G) to simulate real speeches and on a monitor (32-inch display) on the desk in front of them in (T) and (S). To prevent the noise in the environment from interfering with the performance, the audio source of the videos was directly received from the computer for speech-to-text conversion. One video was used in a training session for all participants. We encouraged them to explore every feature and asked them to think about their quick note-taking strategies for GazeNoter based on their experiences in the training session. During the experiment, they watched a TED Talks video in each method and performed note-taking for questions in the Q&A sessions or important points and ideas they wanted to record. After the speeches, they could review the notes, see the transcripts of the notes, or even perform refinement in a 30-second time window, as a small time window between a speech and the Q&A session. Certainly, these could be performed during the speeches.
4.2.3. Measures